# CoAM: Corpus of All-Type Multiword Expressions

### v1

@@ Abstract
Multiword expressions (MWEs) refer to idiomatic sequences of multiple words. MWE identification, i.e., detecting MWEs in text, can play a key role in downstream tasks such as machine translation. Existing datasets for MWE identification are inconsistently annotated, limited to a single type of MWE, or limited in size. To enable reliable and comprehensive evaluation, we created CoAM: Corpus of All-Type Multiword Expressions, a dataset of 1.3K sentences constructed through a multi-step process to enhance data quality consisting of human annotation, human review, and automated consistency checking. MWEs in CoAM are tagged with MWE types, such as Noun and Verb, to enable fine-grained error analysis. Annotations for CoAM were collected using a new interface created with our interface generator, which allows easy and flexible annotation of MWEs in any form, including discontinuous ones.111 The CoAM dataset and source code of the interface generator will be publicly accessible upon acceptance. Through experiments using CoAM, we find that a fine-tuned large language model outperforms the current state-of-the-art approach for MWE identification. Furthermore, analysis using our MWE type tagged data reveals that Verb MWEs are easier than Noun MWEs to identify across approaches.
@@ 1 Introduction
Vocabulary plays a critical role in the comprehension of natural language. Vocabulary is often thought of as consisting of single words, but multiword expressions (MWEs) also form an important part (Jackendoff, 1995). MWEs are defined as idiomatic sequences of multiple words Baldwin and Kim (2010). We focus specifically on those whose meaning or grammatical structure cannot be derived from their constituent words, as they can impede text analysis or comprehension. In other words, we exclude transparent collocations. For example, the bold sequences below constitute MWEs.
@@
MWE identification (MWEI), the process of automatically tagging MWEs in text, has been the object of much research (Constant et al., 2017). MWEI can play a key role in a wide range of downstream tasks, including, but not limited to, machine translation Briakou et al. (2024) and lexical complexity assessment Kochmar et al. (2020). However, as described in Section 2, existing datasets for MWE identification are inconsistently annotated, limited to a single type of MWE, or limited in size. This hinders reliable and comprehensive evaluations of MWEI systems.
In this paper, we introduce the Corpus of All-type Multiword expressions (CoAM), a dataset of 1.3K sentences for MWEI. “Types” refer to MWE categories assigned based on the part of speech of the MWE, with “all-type” signifying the inclusion of all such categories of MWE. To ensure annotation quality, we assigned two annotators and a reviewer to each sentence and checked all annotations for consistency. Additionally, MWEs in the CoAM test set are tagged with their types, such as Noun and Verb, facilitating fine-grained error analysis of MWEI systems. This enables us to address questions such as: “Are verbal MWEs—the focus of the long-running PARSEME project Savary et al. (2017)—more difficult to identify compared to other categories of MWEs?”
@@
Using CoAM, we evaluate two distinct MWEI approaches. The first approach, MWEasWSD (MaW, Tanner and Hoffman, 2023), combines a rule-based pipeline and a trainable bi-encoder model, achieving state-of-the-art performance on the DiMSUM dataset Schneider et al. (2016). The second is LLM fine-tuning for MWEI, inspired by the effectiveness of similar approaches in tasks such as named entity recognition (NER, Zhou et al., 2024). We use CoAM to train and evaluate LLMs from the Llama Dubey et al. (2024) and Qwen Qwen Team (2024) model families.
The results reveal that a fine-tuned Qwen model with 72B parameters greatly outperforms MaW, suggesting the effectiveness of LLM fine-tuning. On the other hand, all approaches suffer from low recall (e.g., 50.7% for Qwen-72B). Further analysis shows that fine-tuned LLMs struggle more to detect Noun and Clause MWEs than to detect Verb MWEs. MWEs not contained in WordNet Miller (1995), e.g., real estate, were found particularly difficult to identify, presumably because they are less widely recognized as MWEs.
@@ 2 Related Work
Previous studies have presented several datasets for MWEI and idiom222 According to Tedeschi et al. (2022), idioms are a subset of MWEs. identification. The DiMSUM Schneider et al. (2016) dataset consists of 5,799 sentences annotated with MWEs, including, but not limited to, verbal MWEs, noun MWEs, phatics, and multi-word (MW) proper nouns. Although DiMSUM has been used in multiple MWEI studies, e.g., Kirilin et al. (2016) and Liu et al. (2021), there exist inconsistencies in the annotation, which hinders proper evaluation. Tanner and Hoffman (2023) found that over 80% of the false positives of their system were actually caused by DiMSUM’s inconsistent annotation. Next, the PARSEME corpus Savary et al. (2017); Walsh et al. (2018) is a high-quality dataset for the identification of verbal MWEs. It has been continuously updated, and the latest version (1.3) contains over 455,000 sentences across 26 languages. However, their focus is limited to verbal MWEs. Lastly, ID10M Tedeschi et al. (2022) is an idiom detection dataset consisting of automatically created training data in 10 languages and a manually curated evaluation benchmark of four languages. Their evaluation dataset was created with the help of professional annotators, but it contains only 200 sentences per language. They used Wiktionary as the source of idioms, and MWEs not in Wiktionary were skipped in the annotation. Additionally, they did not annotate discontinuous idioms.
@@
ID10M used the BIO scheme. DiMSUM employed the more flexible 6-tag scheme, which allows discontinuous MWEs, but cannot handle overlapping MWEs (see Section A.4 for an example of an overlapping MWE). In contrast, the parseme-tsv format Savary et al. (2017) accepts MWEs in any form, which leads us to adopt an equivalent data format.
@@ 3 Task Formulation
We formulate the MWEI task as token-level sequence tagging, where each token can belong to multiple MWEs. The systems are provided with tokenized sentences. Given the sequence of tokens in each sentence, $x_{1},\ldots{},x_{n}$, the task is to output a list of MWEs where each MWE is represented as a list of word indices. The $i$-th MWE in the sentence is represented with $[\mathrm{idx}(t_{1,1}),\mathrm{idx}(t_{1,2}),\ldots{}]$ where $t_{i,j}$ is the $j$-th token of the MWE and $\mathrm{idx}(t)$ is $t$’s identifier, i.e., 1-based index. Our annotation interface allows annotations in this scheme, as illustrated in Figure 1.
@@ 4.1 Data Selection and Preprocessing
In selecting the sources of our dataset, we prioritized sources aimed at general audiences so that they are in standard English and mostly free of grammatical errors. We also included both written and transcripted spoken texts. Consequently, we constructed CoAM from the following four data sources (see Section A.1 for their details). News is news text written by professional writers sourced from EMM NewsBrief, which was introduced by Glavaš and Štajner (2013).333 The dataset also contains data from WikiNews and Wikipedia, but we preferred EMM NewsBrief, which contains more MWEs according to our preliminary analysis based on the annotation by Kochmar et al. (2020). Commentary is commentaries on news from the WMT23 Shared Task monolingual training data (Kocmi et al., 2023). TED is a collection of TED talk transcriptions from two sources. (1) NAIST is the dataset by Neubig et al. (2014); (2) IWSLT is the subset of IWSLT 2017 Shared Task Cettolo et al. (2017). UD is a collection of single sentences sourced from weblogs, reviews, question-answers, newsgroups, and emails in the English Web Treebank. It is part of the Universal Dependencies (Zeman et al., 2018) and the English PARSEME corpus Walsh et al. (2018). We included the UD section of the PARSEME corpus in our dataset to enable better comparison between the annotations. We did not, however, include any sentences from UD in the test set, because UD contains user-generated content with frequent grammatical errors. All the other sources were used both for training and test splits.
For all sources, we took the first 10 (or all, when there are less than 10) sentences from the article or talk and presented them in original order. Each sentence was tokenized using SpaCy’s en_core_web_lg (3.7.1), with the exception of the UD sentences, which were already tokenized. Note that what we call words are tokens given in this manner.
@@ 4.2 Annotation Guidelines
The construction of a reliable dataset requires clear guidelines for annotators in order to minimize misannotation. However, we found no such guidelines for all-type MWE annotation, which led us to create new guidelines based on the MWE definition by Baldwin and Kim (2010) and the PARSEME annotation guidelines.444 https://parsemefr.lis-lab.fr/parseme-st-guidelines/1.3/
We define MWEs as idiomatic sequences that satisfy the following three conditions.
(a) It consists of at least two words that are always realized by the same lexemes. This condition means that his in put yourself in his shoes is not part of the MWE, as other words like Michael’s can replace his.
(b) It displays semantic, lexical, or syntactic idiomaticity. Semantic idiomaticity occurs when the meaning of an expression cannot be derived from its constituent words. It is the most important among idiomaticity types because the majority of MWEs are defined as MWEs due to this type of idiomaticity. See more description on the semantic idiomaticity in relation to non-compositionality in Section A.3. Note that transparent collocations such as stuck at are not an MWE in our definition because they are not semantically, lexically, or syntactically idiomatic. In other words, we did not take statistical idiomaticity as a qualification for an MWE, unlike Baldwin and Kim (2010), for the following reasons. First, our primary aim is to identify sequences that lead to difficulties in reading, and transparent collocations are understandable if the reader knows the constituent words. Second, statistical idiomaticity is difficult for human annotators to judge, and it would bring considerable ambiguity to the definition of MWEs.
(c) It is not a proper noun, i.e., a specific name of a person, facility, and so on. Although some previous studies include MW proper nouns in MWEs Schneider et al. (2016); Baldwin and Kim (2010), we exclude them because the identification of MW proper nouns has already been actively studied in the field of NER. Delegating MW proper noun identification to NER researchers will enable a more focused investigation on our part.
The full MWE definition is in Section A.2. The guidelines were updated whenever an issue was found, such as an ambiguous description.
Figure 1: An example sentence presented in our checkbox-based interface. Checks in the checkboxes are instantly reflected in the bordered zone at the upper right, showing which MWEs are currently marked.
@@ 4.3 Annotation Interface
Tool Flexible annota- tion Simple interface Easy collabora- tion Customiz- able interface
@@
For flexible and efficient annotation, we developed a novel annotation interface generator, CAIGen. CAIGen builds a checkbox-based interface in Google Sheets555 https://developers.google.com/sheets , allowing annotators to perform annotations of MWEs and other kinds of spans simply by checking checkboxes as shown in Figure 1.
Its main advantages over other comparable annotation tools are summarized in Table 1. The first tool, brat666 https://github.com/nlplab/brat Stenetorp et al. (2012), has been used for annotation of a wide range of tasks such as sentiment analysis (Pontiki et al., 2016) and NER (Tabassum et al., 2020). FLAT777 https://github.com/proycon/flat was used for the annotation of the PARSEME corpus Savary et al. (2017). The advantages of CAIGen are as follows. First, the generated interface accepts annotations of spans of any form, including discontinuous or overlapping spans (see Section A.4 for examples), making it more flexible than the previous tools. Second, CAIGen provides a simple spreadsheet-based interface, which is familiar and accessible for technical and non-technical annotators alike. Third, CAIGen alleviates researchers’ overhead of managing servers and annotator accounts by delegating to Google, while the other tools require either (1) distributing data slices to each annotator and having them run applications locally or (2) running and managing the application on a remote server. Lastly, our interface is highly customizable because CAIGen builds it using Google Apps Script, a widely used programming language based on JavaScript; for example, simple customization could allow you to collect annotations about additional information, such as span types.
@@ 4.4 Annotation
The annotation was done by (1) two annotators hired through a company and (2) five authors. The hired annotators comprise one native and one non-native English speaker, both with at least six years of experience as translators. They received feedback from the authors after annotating the first 50 or so sentences to correct misunderstandings about the annotation guidelines. We paid roughly 1 USD per sentence as a reward. Meanwhile, the author annotators consist of three native and two non-native speakers, all with a language-related degree and sufficient English proficiency to perform the task. For reliable annotation, we assigned each sentence to two annotators (one hired annotator and one author), ensuring that at least one was a native English speaker, as the task requires a rich English vocabulary.
@@ 4.5 Inter-Annotator Agreement
With two annotators being assigned to each sentence, we measured the inter-annotator agreement (IAA) by the MWE-based (exact-match) F1 score. The IAA was found to be only 37.3%, which could be attributed to annotators failing to annotate valid MWEs or to subjectivity in idiomaticity judgments despite the carefully written guidelines. We thus suggest that MWE datasets should be constructed with at least two independent annotators.

### v2

@@ Abstract
Multiword expressions (MWEs) refer to idiomatic sequences of multiple words. MWE identification, i.e., detecting MWEs in text, can play a key role in downstream tasks such as machine translation, but existing datasets for the task are inconsistently annotated, limited to a single type of MWE, or limited in size.
To enable reliable and comprehensive evaluation, we created CoAM: Corpus of All-Type Multiword Expressions, a dataset of 1.3K sentences constructed through a multi-step process to enhance data quality consisting of human annotation, human review, and automated consistency checking. Additionally, for the first time in a dataset for MWE identification, CoAM’s MWEs are tagged with MWE types, such as Noun and Verb, enabling fine-grained error analysis.111 The dataset is available at https://huggingface.co/datasets/yusuke196/CoAM. Annotations for CoAM were collected using a new interface created with our interface generator, which allows easy and flexible annotation of MWEs in any form.222 The source code of the interface generator is available at https://github.com/Yusuke196/CAIGen. Through experiments using CoAM, we find that a fine-tuned large language model outperforms MWEasWSD, which achieved the state-of-the-art performance on the DiMSUM dataset. Furthermore, analysis using our MWE type tagged data reveals that Verb MWEs are easier than Noun MWEs to identify across approaches.
@@ 1 Introduction
Vocabulary plays a critical role in the comprehension of natural language. While often simplified as a collection of single words, vocabulary also includes idiomatic sequences known as multiword expressions (MWEs) Baldwin and Kim (2010), which form an important part of language knowledge (Jackendoff, 1995). We focus specifically on sequences whose meaning or grammatical structure cannot be derived from their constituent words, as they can impede text analysis or comprehension; that is, we exclude transparent collocations. For example, the bold sequences below constitute MWEs.
@@
MWE identification (MWEI), the process of automatically tagging MWEs in text (Constant et al., 2017), is valuable for tasks where word-by-word text processing is insufficient, such as machine translation Li et al. (2024) and lexical complexity assessment Kochmar et al. (2020). Briakou et al. (2024) investigated step-by-step translation using a large language model (LLM) and demonstrated that identifying MWEs in the initial step improves the translation quality. Additionally, it is possible to build reading assistance systems to automatically gloss text by combining MWEI with definition modeling (Bevilacqua et al., 2020); as Huang et al. (2022) discusses, definition modeling will help laypeople comprehend specialized terminology, which includes many MWEs.
Despite the importance of MWEI, existing datasets for the task are inconsistently annotated, limited to a single type of MWE, or limited in size, as described in Section 2. This hinders reliable and comprehensive evaluations of MWEI systems.
In this paper, we introduce the Corpus of All-type Multiword expressions (CoAM), a dataset of 1.3K sentences for MWEI. “Types” refer to MWE categories assigned based on the part of speech of the MWE, with “all-type” signifying the inclusion of all such categories of MWEs. To ensure annotation quality, we assigned two annotators and a reviewer to each sentence and checked all annotations for consistency. Additionally, MWEs in the CoAM test set are tagged with their types, such as Noun and Verb, facilitating fine-grained error analysis of MWEI systems. This enables us to address questions such as: “Are verbal MWEs—the focus of the long-running PARSEME project Savary et al. (2017)—more difficult to identify compared to other categories of MWEs?”
@@
Using CoAM, we evaluate two distinct MWEI approaches. The first approach, MWEasWSD (MaW, Tanner and Hoffman, 2023), combines a rule-based pipeline and a trainable bi-encoder model, achieving state-of-the-art performance on the DiMSUM dataset Schneider et al. (2016). The second is LLM fine-tuning for MWEI, inspired by the effectiveness of similar approaches in the task of named entity recognition (NER, Zhou et al., 2024). We use CoAM to train and evaluate LLMs from the Llama Dubey et al. (2024) and Qwen Qwen Team (2024) model families.
The results reveal that a fine-tuned Qwen model with 72B parameters greatly outperforms MaW, demonstrating the effectiveness of LLM fine-tuning. On the other hand, all approaches suffer from low recall (e.g., 50.7% for Qwen-72B). Further analysis shows that fine-tuned LLMs struggle more with detecting Noun and Clause MWEs than with detecting Verb MWEs. MWEs not contained in WordNet Miller (1995), e.g., real estate, were found particularly difficult to identify, presumably because they are less widely recognized as MWEs.
@@ 2 Related Work
Previous studies have presented several datasets for MWEI and idiom333 According to Tedeschi et al. (2022), idioms are a subset of MWEs. identification. The DiMSUM Schneider et al. (2016) dataset consists of 5,799 sentences annotated with MWEs, including, but not limited to, verbal MWEs, noun MWEs, phatics, and multi-word (MW) proper nouns. Although DiMSUM has been used in multiple MWEI studies, e.g., Kirilin et al. (2016) and Liu et al. (2021), there exist inconsistencies in the annotation, which hinder proper evaluation. Tanner and Hoffman (2023) found that over 80% of the false positives of their system were actually caused by DiMSUM’s inconsistent annotation. Next, the PARSEME corpus Savary et al. (2017); Walsh et al. (2018) is a high-quality dataset for the identification of verbal MWEs. It has been continuously updated, and the latest version (1.3) contains over 455,000 sentences across 26 languages. However, their focus is limited to verbal MWEs. Lastly, ID10M Tedeschi et al. (2022) is an idiom detection dataset consisting of automatically created training data in 10 languages and a manually curated evaluation benchmark of four languages. Their evaluation dataset was created with the help of professional annotators, but it contains only 200 sentences per language. They used Wiktionary as the source of idioms, and MWEs not in Wiktionary were skipped in the annotation. Additionally, they did not annotate discontinuous idioms. Among non-English datasets for MWE identification, Sequoia Corpus Candito et al. (2020), a French corpus annotated for MWEs and named entities, stands out for containing 3,440 MWEs.
@@
ID10M uses the BIO scheme. DiMSUM uses the more flexible 6-tag scheme, which allows discontinuous MWEs, but cannot handle overlapping MWEs (see Section D.2 for an example of overlapping MWEs). In contrast, the parseme-tsv format Savary et al. (2017) accepts MWEs in any form, which leads us to adopt an equivalent data format.
@@ 3 Task Formulation
We formulate the MWEI task as token-level sequence tagging, where the inputs are tokenized sentences and each token can belong to multiple MWEs. Given the token sequence of each sentence, $x_{1},\ldots{},x_{n}$, the task is to output a list of MWEs where each MWE is represented as a list of token indices. The $i$-th MWE in the sentence is represented with $[\mathrm{idx}(t_{i,1}),\mathrm{idx}(t_{i,2}),\ldots{}]$ where $t_{i,j}$ is the $j$-th token of the MWE and $\mathrm{idx}(t)$ is $t$’s index. Our annotation interface allows annotations in this scheme, as illustrated in Figure 1.
@@ 4.1 Data Selection and Preprocessing
In selecting the sources of our dataset, we prioritized sources aimed at general audiences so that they are in standard English and mostly free of grammatical errors. We also included both written text and transcribed speech. Consequently, we utilized the following four data sources (see Section A.1 for their details). News is news text written by professional writers sourced from EMM NewsBrief, which was introduced by Glavaš and Štajner (2013).444 The dataset also contains data from WikiNews and Wikipedia, but we preferred EMM NewsBrief, which contains more MWEs according to our preliminary analysis based on the annotation by Kochmar et al. (2020). Commentary is commentaries on news from the WMT23 Shared Task monolingual training data (Kocmi et al., 2023). TED is a collection of TED talk transcriptions from two sources. (1) NAIST is the dataset by Neubig et al. (2014); (2) IWSLT is the subset of IWSLT 2017 Shared Task Cettolo et al. (2017). UD is a collection of single sentences sourced from weblogs, reviews, question-answers, newsgroups, and emails in the English Web Treebank. It is part of the Universal Dependencies (Zeman et al., 2018) and the English PARSEME corpus Walsh et al. (2018). We did not include any sentences from UD in the test set, because UD contains user-generated content with frequent grammatical errors. All the other sources were used both for training and test splits.
For all sources, we took the first 10 (or all, when there are less than 10) sentences from each article or talk and presented them in original order. Each sentence was tokenized using spaCy (we use version 3.7.1 and the en_core_web_trf model throughout this paper), with the exception of the UD sentences, which were already tokenized. Note that what we call words are tokens given in this manner.
@@ 4.2 Annotation Guidelines
The construction of a reliable dataset requires clear guidelines for annotators in order to minimize misannotation. However, we found no such guidelines for all-type MWE annotation, which led us to create new guidelines.
We define MWEs as idiomatic sequences that satisfy the following three conditions, based on the definition by the often-cited Baldwin and Kim (2010) and the PARSEME annotation guidelines.555 https://parsemefr.lis-lab.fr/parseme-st-guidelines/1.3/
(a) It consists of at least two words that are always realized by the same lexemes. This condition means that his in put yourself in his shoes is not part of the MWE, as other words like Michael’s can replace it.
(b) It displays semantic, lexical, or syntactic idiomaticity. Semantic idiomaticity occurs when the meaning of an expression cannot be derived from its constituent words. It is the most important type of idiomaticity because it accounts for the majority of MWEs being classified as idiomatic. See additional discussion on semantic idiomaticity in relation to non-compositionality in Section A.2. Note that transparent collocations such as stuck at are not MWEs in our definition because they are not semantically, lexically, or syntactically idiomatic. The meaning or grammatical structure of transparent collocations can be understood by their constituents, and thus they can be processed word-by-word.
(c) It is not a proper noun, i.e., a specific name of a person, facility, and so on. Previous MWE studies either classified MW proper nouns as MWEs Schneider et al. (2016) or excluded them as non-idiomatic Tayyar Madabushi et al. (2022). We opted for the latter, as proper nouns are linked to encyclopedic knowledge rather than lexicographic knowledge (Navigli and Ponzetto, 2012), to which typical MWEs belong to.
The full MWE definition is given in Appendix D. The guidelines were updated whenever an issue was found, e.g., an ambiguous description.
Figure 1: An example sentence presented in our interface by CAIGen. Checks in the checkboxes are instantly reflected in the bordered zone at the upper right, showing which MWEs are currently marked. Note that the number of rows is reduced here for brevity; the actual interfaces we used have nine rows per sentence.
@@ 4.3 Annotation Interface
For flexible and efficient annotation, we developed a novel annotation interface generator, CAIGen. CAIGen builds a checkbox-based interface in Google Sheets666 https://developers.google.com/sheets , allowing annotators to perform annotations of MWEs and other kinds of spans simply by checking checkboxes as shown in Figure 1. In the interface generation process, for each sentence, CAIGen first writes the sentence ID, the sentence itself, and a bordered box to show the result of annotations. Then, it arranges each token from left to right, wrapping the line when its length hits a pre-set limit.
Flexible annota- tion Easy collabora- tion Easy-to-use interface Customiz- able interface
@@
Its main advantages over other comparable annotation tools are summarized in Table 1. The first tool, brat777 https://github.com/nlplab/brat Stenetorp et al. (2012), has been used for annotation of a wide range of tasks such as sentiment analysis (Pontiki et al., 2016) and NER (Tabassum et al., 2020). FLAT888 https://github.com/proycon/flat was used for the annotation of the PARSEME corpus Savary et al. (2017). The advantages of CAIGen over these are as follows. First, the generated interface supports annotations of spans of any form, including discontinuous or overlapping spans (see Section D.2 for examples), making it more flexible than the previous tools. Second, CAIGen alleviates researchers’ overhead of managing servers and annotator accounts by delegating to Google, while the other tools require either (1) distributing data slices to each annotator and having them run applications locally or (2) running and managing the application on a remote server. Third, CAIGen provides an easy-to-use spreadsheet-based interface, which is simple and familiar to technical and non-technical annotators alike. Lastly, our interface is highly customizable because CAIGen builds it using Google Apps Script, a well-documented programming language based on JavaScript; for example, additional annotations of span types could be easily collected.
@@ 4.4 Annotation
The annotation was done by two annotators hired through a company and five authors. The hired annotators comprise one native and one non-native English speaker, both with at least six years of experience as translators. They received feedback from the authors after annotating the first 50 or so sentences to correct misunderstandings about the guidelines. We paid 154 JPY (roughly 1 USD) per sentence as a reward to the company. Meanwhile, the author annotators consist of three native and two non-native speakers, all with a language-related degree and sufficient English proficiency to perform the task. For reliable annotation, we assigned each sentence to two annotators (one hired annotator and one author), ensuring that at least one was a native English speaker, as the task requires a rich English vocabulary.
@@ 4.5 Inter-Annotator Agreement
With two annotators being assigned to each sentence, we measured the inter-annotator agreement (IAA) by the MWE-based (exact-match) F1 score. The IAA was found to be only 37.3%, which could be attributed to annotators failing to annotate valid MWEs. We thus suggest that MWE datasets should be constructed with at least two independent annotators to reduce missed MWEs.
Nonetheless, it should be noted that this IAA is comparable with other MWEI datasets, and that it does not reflect the quality of CoAM. Our IAA is computed in a strict way, which we chose for its clear interpretability. CoAM’s IAA is comparable to those of other datasets when the same, more lenient methods are used. PARSEME 1.1 Ramisch et al. (2018) computes F1 for exact span matches as we do, but they only report “the highest scores among all possible annotator pairs” (see Table 1, ibid.). The highest score for English PARSEME is 52.9%, similar to the highest score computed in the same way for CoAM: 52.2%. Meanwhile, DiMSUM was annotated by a single annotator, and thus no actual agreement was reported. Although Schneider et al. (2016) estimates DiMSUM’s IAA to range from 60% to 75%, this is based on a very small subset (66 sentences) reannotated by the first author and computed as a more lenient F1 score based on partial span matches. Thus, while our reported IAA value is low, we emphasize that our IAA is in line with comparable MWEI datasets. Additionally, we addressed the annotation disagreements through the combination of human review and consistency checking as described below.
Type PoS of Head Description Example
Verb VERB, AUX Verbal MWEs. stand for, pick up, be the case
Mod/ Conn ADJ, ADV, ADP, CCONJ, SCONJ Adjectival, adverbial, adpositional, or connective MWEs. under the weather, of course, in spite of
Clause VERB, AUX MWEs containing (1) a verb or auxiliary verb and (2) its nominal subject, such as phatics and proverbs. you’re welcome, the early bird gets the worm, when it comes to
Other Any MWEs whose head is not contained in them or whose PoS is none of the PoS above. and so on, oh man